27-Y^2=0
and
square root of9x/2 - 2 +7

Answer by stanbon(75887)

27-Y^2=0
Factor to get:
(3sqr3-y)(3sq3+y) = 0
y=3sqrt3
y=-3sqrt3
------------------
and
square root of9x/2 - 2 = 7
sqrt(9x/2) -2 = 7/2
sqrt(9x/2) = 9
Square both sides to get:
9x/2 = 81
9x = 2*81
x = 2*9 = 18
===============
Cheers,
Stan H.